Explain in your own words what the meaning of domain is. Also, explain why a denominator cannot be zero.

Find the domain for each of the two rational expressions below.

Write the domain of each rational expression in set notation (as demonstrated in the example).

Do both of your rational expressions have excluded values in their domains? If yes, explain why they are to be excluded from the domains. If no, explain why no exclusions are necessary.

Incorporate the following five math vocabulary words into your discussion. Use bold font to emphasize the words in your writing. Do not write definitions for the words; use them appropriately in sentences describing your math work.

Domain

Excluded value

Set

Factor

Real numbers

Your initial post should be at least 250 words in length. Support your claims with examples from required material(s) and/or other scholarly resources, and properly cite any references.

16b2 – 1
2b2 + 11b – 6

c2 + 2c – 10
10c2 – 80c + 160
Answer #1

\\\frac{16b^2-1}{2b^2+11b-6}\\ \\\frac{(4b-1)(4b+1)}{2b^2+12b-b-6}\\ \\\frac{(4b-1)(4b+1)}{(2b-1)(b+6)}\\

Domain - value which can be inputed in this rational expression. R - {1/2, -6}

on the value 1/2 and -6, the expression becomes undefined.

Excluded value = 1/2 and - 6

Set = (-infitnity, -6) U (-6, 1/2) U (1/2, infinity)

Factor = (4b-1) (4b+1)

As per Chegg rules, you can get answers to just 1 question (4 parts) in one post. Please post the other questions in a separate post.Please upvote the answer.